Outback out r ters sells recreational equipment. One of the companys products, a small camp stove, sells for $1 40 per unit Variable expenses are串98 per stove, anc fixed expenses associated with the stove tota $197,400 per month. Required: 1. Compute the companys break-even point in unit sales and in dollar sales. Polnt Number of stowes Total sales dollars 4,70 658,000 $ 2. If the variable expenses per stove increase as a percentage of the selling price, will it result in a higher or a lower break-even point?(Assume that the fixed expenses remain unchanged.) Higher break-veri poin Lower break even point 3 At present, the company is selling 19,00 stoves per month. The sales mana er is oonvinced that a 10% reduction in the selling price would result in a 259 income statements, one under present operating conditions, and one as operations would appear after the proposed changes increase in monthly sales of st es. Prepare two contribution ormat Present 19,000 Stoves Stoves Total Per unit Total Per unit 5 2650,00 140г$ 2,992,500 120 riable expenses 08 2,650,000S 12 2,992,500 S 197,400 Contribution margin ixcd cxpcriscs 197.400 t operatng income 5 2,452,600 S 2,795,1002·itthe variable expenses per stove increase as a percentage of the selling price, will it result in a higher or a lower break-even point? (Assume that the fixed expenses remain unchanged.) Higher break-even point 令Lower break-even point 3. A present, he company is selling 19 000 sto es per month. The sales manac er is convinced that a 10% reduction in the selling p ce would result in a 25% increase in monthly sales of stoves. Prepare two contribution or at income statements, one under present operating conditions, and one as operations would appear after the proposed changes. Outback Outfitters Present 19,000 Stoves Stoves Total Per unit Total Per unit 140「$ 2,992.5001 s 128 riable expenses 98 89 Contribistion margin 42 ixed expenses 197 400 197,400 t operatng income 5 2452,600 S 2,795,100 4. Refer to the data in (3) above. How many stoves would have to be sold at the new selling price to yield a minimum net operating income of 577,000 per month? (Round your answer to the nearest whole number.)
Answer #1
1
Break-even point
Number of stoves 4700 =197400/(140-98)
Total sales dollars 658000 =4700*140
2
Higher break even point
3
   Present Proposed
19000 stoves 23750 stoves
Total Per unit Total Per unit
Sales 2660000 140 2992500 126
Variable expenses 1862000 98 2327500 98
Contribution margin 798000 42 665000 28
Fixed expenses 197400 197400
Net operating income 600600 467600
4
Number of stoves to be sold 9800 =(197400+77000)/28